OpenStudy (anonymous):

Please help me with this geometry problem, I really don't understand and its making me frustrated. Thank you so much Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area. How much area will each delivery radius cover? Write the equation for each circle created. Please show your work for all calculations.

OpenStudy (anonymous):

What's the difficulty exactly?

OpenStudy (anonymous):

The first part is this Determine how many slices you will cut your pizza into. You need to pick two different numbers of slices. Once you have determined the number of slices, calculate the following for both ways of cutting your pizza: the interior angle and arc length of the slices the area of each slice So I chose 8 and 10 for my two slices Than I divided 360/8 which is 45 and 360/10 which is 36

OpenStudy (anonymous):

Um..OK but that has nothing to do with the question you posted?

OpenStudy (anonymous):

the question I posted is asking me to create a equation so I don't get it

OpenStudy (anonymous):

Write the equation for each circle created. This is the troublesome part?

OpenStudy (anonymous):

For a circle centred at (0,0) in the Cartesian plane, the general form of equation is x^2+y^2 = r^2 where r is the radius So if you chose 2 miles and 4 miles for your delivery radii then you would have 2 equations x^2+y^2 = 4 and x^2+y^2 = 16 The area for these circles is just pi*r^2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

ur welcome.